OpenStudy (anonymous):

What is the solution of the system? Use substitution. y = x + 6 y = 2x

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

6 and what

OpenStudy (anonymous):

y = x + 6 y = 2x If y = 2x, then x = 6. y = 2x or y = 2(6) or y = 12 x + 6 or 6 + 6 So… 12 = 6 + 6… Correct?

OpenStudy (anonymous):

Do you understand, now? Or do you need more help?

OpenStudy (anonymous):

i understand thank you.